### v1

@@ 1 Introduction
The self-attention mechanism in ViTs enables more accurate classification by considering the entire image context, which enhances accuracy and robustness compared to traditional CNNs [24]. For example, Dosovitskiy et al. demonstrated that ViTs could outperform CNNs in image classification tasks, highlighting their potential for diversified applications [[22]]. Recent studies have also applied ViTs to plant disease identification [25, 26, 27]. In this study, we applied ViTs to drought stress identification to (a) achieve higher accuracy in detecting drought stress, b) leverage the ViT’s superior ability to handle complex patterns and variations inherent in field images captured in natural settings, ensuring the model’s practicality and relevance to real-world conditions, c) interprete model by utilizing ViTs to generate attention maps, which provide insights into the model’s decision-making process and contribute to the transparency and explainability of the predictions.
@@ 2.1 Preparing the Data
The potato crop aerial images utilized in this study have been sourced from a publicly accessible dataset that encompasses multiple modalities [28]. Collected from a field at the Aberdeen Research and Extension Center, University of Idaho, these images serve as valuable resources for training machine learning models dedicated to crop health assessment in precision agriculture applications. Acquired using a Parrot Sequoia multi-spectral camera mounted on a 3DR Solo drone, the dataset features an RGB sensor with a resolution of $4,608\times 3,456$ pixels and four monochrome sensors capturing narrow bands of light wavelengths: green (550nm), red (660nm), red-edge (735nm), and near-infrared (790nm), each with a resolution of $1,280\times 960$ pixels. The drone flew over the potato field at a low altitude of 3 meters, with the primary objective of capturing drought stress in Russet Burbank potato plants attributed to premature plant senescence.
@@
The dataset consists of 360 RGB image patches in JPG format, each measuring 750×750 pixels. These patches were derived from high-resolution aerial images through cropping, rotating, and resizing operations. The dataset is divided into a training subset of 300 images and a testing subset of 60 images. Ground-truth annotations, provided in both XML and CSV formats, indicate regions of healthy and stressed plants marked by rectangular bounding boxes. These annotations were manually created using the LabelImg software. An example RGB image and its annotated regions are shown in Fig. 1. The testing subset is independent of the training subset and comes from different aerial images. Additionally, the dataset includes corresponding image patches from spectral sensors with red, green, red-edge, and near-infrared bands, each sized 416×416 pixels. However, we are only utilizing the RGB images due to the limitations of the low-resolution monochromatic images.
We utilized the augmented dataset prepared by Butte et al., expanding the original set from 300 to 1,500 images for training. The test set comprised 60 images. From each augmented and test image, annotated windows (i.e., rectangular bounding boxes) were extracted, creating separate folders for ”healthy” and ”stressed” images. These folders served as input for our proposed model. The final count of training images for the ”stressed” and ”healthy” classes is 11,915 and 8,200, respectively. The test set, consisting of 60 images, yielded 401 healthy and 734 stressed images. We evaluated our method on this test set and further assessed the model performance through 5-fold cross-validation.
@@ 2.3 ViT with Transfer Learning
The proposed framework, as shown in Fig 2(b) effectively combines transfer learning, the power of the Vision Transformer and attention-based interpretability to address the challenging task of drought stress identification in potato crop images captured in natural settings. A core component of this approach is the utilization of pre-trained weights. This technique, known as transfer learning, involves leveraging knowledge gained from solving one problem (often a large-scale image classification task) and applying it to a different but related problem. By employing pre-trained weights, the model can benefit from the rich feature representations learned from a massive dataset, accelerating training and potentially improving performance. Using this the ViT is trained over multiple epochs using a combination of binary cross-entropy loss, Adam optimizer and learning rate tuning. Experimenting with different learning rates is essential to find the optimal value for convergence and generalization. Eventually, the classification layer is responsible for making the final prediction. It takes the output of the ViT model and maps it to two classes: ”healthy” and ”stressed.” This layer typically consists of a fully connected neural network with a sigmoid activation function for binary classification.
@@ 3 Results and Discussion
In this section, we present the experimental results of our proposed model for identifying drought stress in potato crop field images. First, we distinguish between healthy and stressed images. Then, we identify the spatial features responsible for the stress. Our experiments with the proposed Vision Transformer (ViT) framework were conducted in three ways:
  • ViT with Transfer Learning: Leveraging pre-trained weights.
  • ViT+SVM with Pre-trained Weights.
@@
The PyTorch library and it’s various sub-packages are leveraged for deep learning functionalities. In addition to this, sklearn, pandas, numpy packages are used for developing, training, and evaluating deep learning models, as well as for performing data analysis and pre-processing. The data pre-processing involves resizing images to (224, 224) pixels and transforming them into tensors.
@@ 3.1 Performance of ViT with Transfer Learning
We relied on PyTorch, a core library for deep learning, which provides the essential tools for tensor computations and neural network construction. Specifically, torch and torch.nn were used to build and train the neural network. To handle image data, we used Torchvision’s models for accessing pre-trained architectures and transforms for preprocessing images. Additionally, TQDM was employed to create progress bars, enhancing visibility into the training process. For data analysis and manipulation, we leveraged scikit-learn for machine learning utilities and Pandas for managing structured data.
The dataset preparation involved mapping class names to numerical labels using class_mapping, where ’healthy’ images were assigned the label 0 and ’stressed’ images were assigned the label 1. We applied image transformations to resize the images to 224x224 pixels and convert them into tensor format, ensuring compatibility with the ViT model. The datasets were loaded using ImageFolder from Torchvision, which facilitated the organization of images into training and testing sets. Data loaders were then created to handle batching, shuffling, and efficient data retrieval during training and validation, with a batch size of 128.
@@
To adapt the Vision Transformer (ViT) architecture (as depicted in Fig 2(a)) for our specific task of binary classification, we began by configuring the model using the models.vit_b_16 variant, setting pretrained=False to start with a clean slate. This approach allowed us to build the model from scratch without any influence from pre-existing learned features. Subsequently, we loaded custom pre-trained weights into the ViT model to realize the approach as shown in Fig 2(b). This step was crucial as it transferred learned representations from a previously trained model to our current architecture, leveraging prior knowledge to enhance performance. A custom class was defined to extend the ViT model by adding a fully connected layer for binary classification. This class also included methods to capture attention weights, which are essential for understanding the model’s focus on different image regions.
The training process involved setting up a binary cross-entropy loss function to evaluate the performance of the classification model. We used the Adam optimizer with a learning rate of 0.001 and 0.0009 to update the model’s weights. The training loop consisted of multiple epochs, during which the model was trained on the training dataset and validated on the test dataset. Training and validation losses were computed to assess the model’s performance and adjust parameters accordingly. Throughout the process, attention weights were captured to provide insights into the model’s decision-making process.
Figure 3: Trainig Losses ViT from scratch vs ViT with Pretrained Weights
(a) Epoch 1 to 20: lr=0.001, Attention dropout=0, MLP dropout=0
(b) Epoch 1 to 10: lr=0.001, Epochs 11 to 20: lr=0.0009, In both cases, Attention dropout=0, MLP dropout=0
(c) Epoch 1 to 8: lr=0.001, Attention dropout=0, MLP dropout=0; Epochs 9 to 20: lr=0.001, Attention dropout=0.2, MLP dropout=0.1
(d) Epoch 1 to 8: lr=0.001, Attention dropout=0, MLP dropout=0; Epochs 9 to 20: lr=0.0009, Attention dropout=0.2, MLP dropout=0.1
Figure 4: Training and Validation Losses
The learning curve typically plots both the training loss and validation loss curves on the same graph. This allows to analyze how the model is performing on both the data it’s trained on (training loss) and unseen data (validation loss). The plot shown in Fig 3 compares the training losses over the first 10 epochs for VIT trained from scratch versus VIT trained with pre-trained weights. The Vision Transformer (ViT) with pre-trained weights demonstrates a notably faster and more consistent reduction in loss compared to the ViT trained from scratch. This indicates superior initial performance and quicker convergence. As a result, we chose to further experiment with pre-trained weights rather than training the model from scratch.
The training loss is plotted in Fig 4(a) for 20 epochs when using a Vision Transformer (ViT) with pre-trained weights and a learning rate of 0.001. The goal of training a machine learning model is to minimize the loss on the validation set. In an ideal scenario, the training loss and validation loss will both decrease as the model trains. However, this is not always the case. In this graph, the training loss decreases steadily over the course of training. This suggests that the model is learning the training data well. However, the validation loss starts to increase after a certain number of epochs, specifically after epoch 8. This is a sign of over-fitting. Over-fitting is a problem that occurs when a model learns the training data too well, and it is no longer able to generalize to unseen data.
The training loss is plotted in Fig 4(b) for 20 epochs using a Vision Transformer (ViT) with pre-trained weights. The learning rate is 0.001 for epochs 1 to 10 and 0.0009 for epochs 11 to 20. The training loss decreases steadily, indicating the model is learning the training data well. However, the validation loss fluctuates and increases after epoch 6, suggesting over-fitting. The training loss is plotted in Fig 4(c) for 20 epochs using a Vision Transformer (ViT) with pre-trained weights. The learning rate is 0.001 for epochs 1 to 8, with no dropout. For epochs 9 to 20, the learning rate remains 0.001, with an attention dropout of 0.2 and an MLP dropout of 0.1. The training loss decreases steadily, but the validation loss fluctuates and increases after epoch 8, indicating over-fitting once dropout is applied. The training loss is plotted in Fig 4(d) for 20 epochs using a Vision Transformer (ViT) with pre-trained weights. The learning rate is 0.001 for epochs 1 to 8 and 0.0009 for epochs 9 to 20. An attention dropout of 0.2 and an MLP dropout of 0.1 are applied for epochs 9 to 20. The training loss decreases steadily, but the validation loss fluctuates and increases after epoch 8, suggesting over-fitting as dropout is applied and the learning rate is adjusted.
In summary, all figures show a decrease in training loss, indicating the model is learning from the training data. Validation loss generally starts to increase or plateau after a certain number of epochs, suggesting overfitting is a common issue in these experiments. The introduction of dropout in Figures 4(c) and 4(d) seems to have exacerbated the over-fitting problem. However, the optimal result for the specified test set, reaching an accuracy of 0.929, is attained with a learning rate of 0.001 and without applying attention dropout or MLP dropout. The Vision Transformer (ViT) with transfer learning demonstrated impressive performance, achieving a precision of 0.95 and recall of 0.94 for stressed images, as well as a precision of 0.89 and recall of 0.91 for healthy images. The F1-scores for stressed and healthy images were 0.95 and 0.90, respectively, highlighting the model’s strong capability in accurately classifying both categories.

### v2

@@ 1 Introduction
The self-attention mechanism in ViTs enables more accurate classification by considering the entire image context, which enhances accuracy and robustness compared to traditional CNNs [24]. For example, Dosovitskiy et al. demonstrated that ViTs could outperform CNNs in image classification tasks, highlighting their potential for diversified applications [22]. Recent studies have effectively applied ViTs both in customized form [25, 26] and CNN- ViT hybrid form to plant disease identification [27]. Thakur et al. [28] developed a lightweight model that combines convolutional blocks from VGG 16 and Inception V7 with transformer components such as multi-head attention and multi-layer perceptron to effectively identify a wide range of plant diseases across multiple crops. The model leverages the local feature extraction capabilities of CNNs and the global feature modeling strength of vision transformers, enabling simultaneous extraction of both local and global features from images. ViT has outperformed Inception V3 in terms of accuracy when distinguishing among nine different tomato leaf disease classes [29]. Parez et al. fine-tuned the vision transformer with a four-fold reduction in training parameters, achieving higher accuracy compared to CNN models for disease identification across three datasets [30]. Yu et al. [31] proposed a framework that integrates soft split token embedding and depth-wise convolutional modules into the vision transformer architecture, resulting in improved accuracy. Replacing the MLP module in the ViT encoder block with an Inception module improved accuracy in multi-crop disease classification while reducing the number of trainable parameters [32]. Thai et al. optimized the ViT architecture for cassava leaf disease detection by pruning less important attention heads and using sparse matrix operations, achieving a 2% improvement in F1-score along with reduced model size and training costs [33]. Hemalatha et al. developed a plant disease localization and classification model which uses co-scale, co-attention, and cross-attention mechanisms with a vision transformer in a multi-task learning framework [34]. Li et al. integrated a convolutional block attention module into the standard ViT encoder, enabling the network to filter out irrelevant information and focus on essential features, leading to improved crop disease classification in rice, wheat, and coffee [35]. Vallabhajosyula et al. [36] proposed a novel framework that combines a transformer encoder with ResNet9 for plant disease classification, outperforming several classical CNN-based models. From the extensive literature review, we observe that while most existing works focus primarily on improving accuracy, few address the reduction of trainable parameters, and none explore the explainability of transformers. Our proposed work differs in two key aspects: first, by deciphering the attention mechanism, we emphasize model explainability; second, we devised ViT and support vector machine (SVM) combined framework and investigated its performance.
In this study, we devised a Vision Transformer (ViT)-based framework and fine-tuned it specifically for drought stress identification, achieving improved accuracy in detecting stress conditions. To demonstrate the model’s interpretability, we employed the inherent self-attention properties of Vision Transformers to produce attention maps. These maps offer meaningful insights into how the model arrives at its decisions, thereby increasing both the transparency and trustworthiness of its predictions. Additionally, we proposed a hybrid ViT+SVM framework that combines the rich feature representation capabilities of ViTs with the strong classification performance of Support Vector Machines (SVMs), resulting in a more robust drought stress identification model.
@@ 2.1 Preparing the Data
The potato crop aerial images utilized in this study have been sourced from a publicly accessible dataset that encompasses multiple modalities [37, 17]. Collected from a field at the Aberdeen Research and Extension Center, University of Idaho, these images serve as valuable resources for training machine learning models dedicated to crop health assessment in precision agriculture applications. Acquired using a Parrot Sequoia multi-spectral camera mounted on a 3DR Solo drone, the dataset features an RGB sensor with a resolution of $4,608\times 3,456$ pixels and four monochrome sensors capturing narrow bands of light wavelengths: green (550nm), red (660nm), red-edge (735nm), and near-infrared (790nm), each with a resolution of $1,280\times 960$ pixels. The drone flew over the potato field at a low altitude of 3 meters, with the primary objective of capturing drought stress in Russet Burbank potato plants attributed to premature plant senescence.
@@
The dataset comprises of 360 RGB image patches in JPG format, each measuring 750×750 pixels. These patches were derived from high-resolution aerial images through cropping, rotating, and resizing operations. Data augmentation was applied to an initial set of 300 images as per procedure in Butte et al. [17], expanding the training dataset to 1,500 images. The remaining 60 images were reserved exclusively for testing. No data augmentation was performed on the test images to ensure an unbiased evaluation. Training classification models requires labeled data with annotated regions of interest. In this study, the targets were regions containing healthy and stressed potato plants. These two conditions were visually distinguishable based on color—healthy plants appeared green, whereas stressed plants exhibited a yellowish hue. Manual annotation was performed using the open-source graphical tool LabelImg [38], allowing bounding boxes to be drawn around both healthy and stressed regions. The resulting annotations, including class labels and bounding box coordinates, were saved and used to generate ground truth data for training the proposed models.
Additionally, the dataset includes corresponding image patches from spectral sensors with red, green, red-edge, and near-infrared bands, each sized 416×416 pixels. However, we are only utilizing the RGB images due to the limitations of the low-resolution monochromatic images.
The augmented dataset, consisting of 1,500 images, was used for training, while the test set included 60 distinct images. From both training and test images, annotated windows (i.e., rectangular patches) were extracted based on the bounding box annotations. Each extracted window was labeled as either “healthy” or “stressed”, corresponding to the visual condition of the crop. As illustrated in Fig 1, the original image is shown in Fig 1(a), and the corresponding extracted windows are depicted in Fig 1(b). In this example, six windows represent healthy regions, while three represent stressed regions. Following this extraction process, the final training dataset comprised 11,915 stressed and 8,200 healthy image patches. From the 60 test images, we obtained 734 stressed and 401 healthy windows. The performance of the proposed model was first evaluated on this test set, and further validated through 5-fold cross-validation to ensure robustness and generalization.
@@ 2.3 ViT with Transfer Learning
The proposed framework, as shown in Fig 2(b) effectively combines transfer learning, the power of the Vision Transformer and attention-based interpretability to address the challenging task of drought stress identification in potato crop images captured in natural settings. A core component of this approach is the utilization of pre-trained weights. This technique, known as transfer learning, involves leveraging knowledge gained from solving one problem (often a large-scale image classification task) and applying it to a different but related problem. Specifically, we used the model initialized with pre-trained weights from training on the ImageNet-1k dataset with 1000 classes. By employing pre-trained weights, the model can benefit from the rich feature representations learned from a massive dataset, accelerating training and potentially improving performance. Using this the ViT is trained and fine-tuned over multiple epochs using a combination of binary cross-entropy loss, Adam/AdamW optimizer and learning rate tuning. Experimenting with different learning rates is essential to find the optimal value for convergence and generalization. Eventually, the classification layer is responsible for making the final prediction. It takes the output of the ViT model and maps it to two classes: ”healthy” and ”stressed.” This layer typically consists of a fully connected neural network with a sigmoid activation function for binary classification.
@@ 3 Results and Discussion
In this section, we present the experimental results of our proposed model for identifying drought stress in potato crop field images. First, we distinguish between healthy and stressed images. Then, we identify the spatial features responsible for the stress. Our experiments with the proposed Vision Transformer (ViT) framework were conducted in two ways:
  • ViT with Transfer Learning (ViT-TL): Leveraging pre-trained weights.
@@
We used the PyTorch library and its sub-packages to implement deep learning functionalities, particularly employing torch and torch.nn for tensor operations and neural network construction. For handling image data, we utilized Torchvision’s models for accessing pre-trained architectures and transforms for preprocessing, which included resizing images to (224, 224) pixels and converting them into tensors. Additionally, TQDM was used to generate progress bars for better training visibility. For data analysis and pre-processing, we employed pandas, NumPy, and scikit-learn for structured data manipulation, numerical computations, and machine learning utilities.
@@ 3.1 Performance of ViT with Transfer Learning
To adapt the Vision Transformer (ViT) architecture (as depicted in Fig 2(a)) for our specific task of binary classification, we began by configuring the model using the models.ViT-B/16 variant. Subsequently, we loaded custom pre-trained weights into the ViT model to realize vision transformer with transfer learning approach as shown in Fig 2(b). This step was crucial as it transferred learned representations from a previously trained model to our current architecture, leveraging prior knowledge to enhance performance. A custom class was designed (as shown in Algorithm 1) to configure encoder layers as trainable or frozen, with methods to adjust various parameters. It also includes functionality to capture attention weights, crucial for analyzing the model’s focus on specific image regions.
Table 1: Training parameters of the model under different scenarios.
Scenario Model No. of Trainable Layers Learning Rate & Optimizer Callback Parameters Batch Size Attention Dropout MLP Dropout
Patience Factor
Scenario 1 ViT-B/16 Last encoder block 0.001(Adam) 5 0.2 128 0 0
Scenario 2 ViT-B/16 Last two encoder blocks 0.001(Adam) 5 0.2 128 0 0
Scenario 3 ViT-B/16 Last three encoder blocks 0.001(AdamW) 5 0.2 128 0 0
Scenario 4 ViT-L/16 Last three encoder blocks 0.001(AdamW) 5 0.2 128 0 0
Scenario 5 ViT-B/16 Last three encoder blocks 0.001(AdamW) 2 0.2 128 0 0
Scenario 6 ViT-B/16 Last three encoder blocks 0.001(AdamW) 2 0.2 64 0 0
Scenario 7 ViT-B/16 Last two encoder blocks 0.001(AdamW) 5 0.2 128 0.1 0.1
Scenario 8 ViT-B/16 Last two encoder blocks 0.001(AdamW) 5 0.2 128 0.1 0.2
Scenario 9 ViT-B/16 All encoder blocks 0.001(AdamW) 5 0.2 128 0 0
Scenario 10 ViT-B/16 All encoder blocks 0.001(AdamW) 5 0.2 64 0 0
Scenario 11 ViT-B/16 Last two encoder blocks 0.001(Adam) 5 0.2 128 0.1 0.2
Table 1 presents the training configurations adopted across eleven experimental scenarios involving Vision Transformer (ViT)-based models. To balance accuracy with computational efficiency, we primarily fine-tuned a limited number of encoder layers within the pre-trained ViT models instead of retraining the entire architecture. Specifically, the number of trainable encoder blocks was varied across scenarios, starting with only the final block in Scenario 1 and gradually increasing up to all encoder blocks in Scenarios 9 and 10. The ViT-B/16 model, known for its relatively lightweight design, served as the backbone for most experiments. An exception was Scenario 4, where we employed the larger ViT-L/16 model, which has approximately four times the number of parameters compared to ViT-B/16. Despite its increased capacity, ViT-L/16 did not yield a notable improvement in accuracy, leading us to retain ViT-B/16 in subsequent scenarios for better scalability and efficiency. Key training elements were systematically varied to evaluate their effects. Both Adam and AdamW optimizers were tested with a fixed learning rate of 0.001, with AdamW being preferred in most cases due to its improved regularization capabilities. Callback settings included early stopping and learning rate reduction on plateau, controlled by a patience of 5 and a factor of 0.2. However, Scenarios 5 and 6 employed a reduced patience of 2 to accelerate convergence. Batch sizes were set to either 64 or 128 to investigate their influence on model convergence and generalization. Finally, to combat overfitting, additional attention and MLP dropout layers were integrated in Scenarios 7, 8, and 11.
Table 2: Model performance across different scenarios.
Scenario Training Accuracy Validation Accuracy Training Loss Validation Loss Test Accuracy Epoch No.
Scenario 1 0.9899 0.9816 0.0272 0.0466 0.9039 16
Scenario 2 0.9929 0.9831 0.0190 0.0563 0.9083 20
Scenario 3 0.9939 0.9906 0.0178 0.0294 0.9057 16
Scenario 4 0.9431 0.9443 0.1468 0.1433 0.8960 14
Scenario 5 0.9935 0.9901 0.0191 0.0322 0.8819 17
Scenario 6 0.9940 0.9876 0.0180 0.0370 0.8995 16
Scenario 7 0.9833 0.9796 0.0421 0.0444 0.9127 18
Scenario 8 0.9765 0.9747 0.0613 0.0876 0.9162 20
Scenario 9 0.9377 0.9513 0.1589 0.1344 0.9119 16
Scenario 10 0.9320 0.9274 0.1721 0.1747 0.9075 19
Scenario 11 0.9707 0.9672 0.0780 0.0896 0.8942 16
Table 2 summarizes the training and validation accuracy and loss for each scenario along with the final test accuracy and the epoch number at which early stopping was triggered. Across most scenarios, the training and validation accuracies exceed 97%, demonstrating strong convergence. Scenario 5, despite high training accuracy (99.35%), showed comparatively lower test accuracy (88.19%), indicating potential overfitting. In contrast, Scenario 8 exhibited the highest generalization with a test accuracy of 91.62%.
Table 3: Confusion matrix components and test accuracy across different scenarios.
Scenario TP TN FP FN Test Accuracy
Scenario 1 647 379 22 87 0.9039
Scenario 2 661 370 31 73 0.9083
Scenario 3 650 378 23 84 0.9057
Scenario 4 638 379 22 96 0.8960
Scenario 5 617 384 17 117 0.8819
Scenario 6 639 382 19 95 0.8995
Scenario 7 663 373 28 71 0.9127
Scenario 8 661 379 22 73 0.9162
Scenario 9 645 390 11 89 0.9119
Scenario 10 638 392 9 96 0.9075
Scenario 11 637 378 23 97 0.8942
The confusion matrix components for each scenario are provided in Table 3, including the number of true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN). These values corroborate the overall test accuracy and provide insights into class-wise prediction reliability. Scenario 8 again stands out with a balanced count of TP and TN and a lower FN, contributing to its highest accuracy.
Figure 3: Loss curves for 11 scenarios: Fig. a–k corresponding to scenario 1 to 11.
Figure 4: Accuracy curves for 11 scenarios: Fig. a–k corresponding to scenario 1 to 11.
The loss and accuracy curves for all 11 scenarios are depicted in Fig. 3 and Fig. 4, respectively. These plots offer visual confirmation of training convergence and generalization. From the loss and accuracy curves, it is evident that most scenarios converge smoothly, with minimal overfitting. Scenario 5 is a notable exception, exhibiting a pronounced gap between training and validation performance. This is further reflected in its confusion matrix, where the number of false negatives (117) significantly exceeds other scenarios, explaining its poor generalization (test accuracy: 88.19%). Scenario 8, despite modest fluctuations during training, achieves the highest test accuracy (91.62%) and maintains a balanced distribution of true and false predictions. Scenarios 9 and 10 notably report the lowest false positive counts (11 and 9, respectively), making them suitable for applications where false alarms are critical. In contrast, Scenario 7 shows the highest recall (TP: 663, FN: 71), which is necessary in settings where missing stressed cases could be detrimental.
This systematic experimentation helped us analyze the trade-off between model complexity and performance. The results showed that fine-tuning only the last 2–3 encoder blocks of ViT-B/16, combined with suitable regularization, yielded performance comparable to that of training the entire network or using ViT-L/16, but with significantly fewer trainable parameters and faster training times.